OpenStudy (anonymous):

If y varies INVERSELY as x and y=6 when x=18, then find x when y=10.

OpenStudy (dumbcow):

\[y =\frac{k}{x}\] Use x=18 y=6 to solve for k

OpenStudy (anonymous):

since xy = constant , say c c = 18 * 6 so y = 10 => x = 108/10 = 10.8